Shift report 20010910 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Sep 10 2001 - 00:06:13 EDT

  • Next message: Borge Svane Nielsen: "Re: Optical Algorithm for RHIC PID"

    ------------------------------------------------------------
    Date: Mon Sep 10 00:06:13 2001, Shift: 16:00-24:00
    Supervisor: EJ, with: 
    Log book#8 pages 101 - 103
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam until 21.15. Then again from 21.56 and still
       running.
    2. Runs with 1/5 field on FS magnets:
       #4654 : 111K trig 1 events
       #4655 : 128K trig 1 events 
       #4656 : 104K trig 1 events
       #4657 :   6K trig 1 events
               ----
               349k trig 1 events
       Runs with 1/3 field on FS magnets:
       #4661 : 318K trig 1 events
       On all runs D5 were at polarity B and 500 Amps.
    3. T5 tripped in run 4654 and again when beam was 
       restored at 21.58. No other problems. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Sep 10 2001 - 00:06:33 EDT